OpenStudy (anonymous):

I have trouble starting these :/ f(x) = 2x + 7 and f(x) = -1

OpenStudy (anonymous):

Plug the value of f(x), which is -1, into the function. 2(-1)+7= -2+7=5

OpenStudy (anonymous):

I am not sure what the question is? are they asking to find x ? @sujuexob2st that would be if they gave f(-1) and f(x) = 2x + 7 :)

OpenStudy (anonymous):

@Emz.Yaoi oh true, I never though of that, but I just assumed that it would be the case of f(-1)

OpenStudy (anonymous):

its coming straight off my homeqwork ... but i think 5 is right

OpenStudy (anonymous):

I was guessing f(x) = f(x) so 2x + 7 = -1 solve for x

OpenStudy (anonymous):

if which case it would be -4, not 5

OpenStudy (anonymous):

Well its multipal choice and both awnsers are on there ..

OpenStudy (anonymous):

f(2) if f(x) = (x + 1)2 what about this one ?

jigglypuff314 (jigglypuff314):

that would would be f(2) means to plug in 2 for x :)

OpenStudy (anonymous):

Thank you !!

jigglypuff314 (jigglypuff314):

you're welcome :)